OpenStudy (anonymous):

plzzzz help!!! –(9 + (–17)) = –9 + x

OpenStudy (anonymous):

what does need x equal to make this eqation true

OpenStudy (anonymous):

@Goose.

OpenStudy (melodycrazysouthern):

i can halp

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

does X maybe equal -17?

OpenStudy (melodycrazysouthern):

x-9=17

OpenStudy (anonymous):

so x=8?

OpenStudy (anonymous):

@melodycrazysouthern ??

OpenStudy (melodycrazysouthern):

1x - 9 = 17 1x = 17 + 9 1x = 26 (divide both sides by 1 to get x) 1x/1 = 26/1 x = 26

OpenStudy (anonymous):

oh

OpenStudy (melodycrazysouthern):

im in 8th grade

OpenStudy (anonymous):

ah!